Let fn(x)= n/(x+2n) is uniformly convergent in [0,k] , k>0

1
Expert's answer
2022-03-16T13:06:43-0400

ANSWER Let "x\in[0,k]" , then "\lim_{x\rightarrow\infty}\frac{n}{x+2n}=\lim_{x\rightarrow\infty}\frac{1}{\frac{x}{n}+2 }=\frac{1}{2}" .

So, the sequence converges pointwise ( "f_{n}\rightarrow f" ) on "[0,k]" ,where "f(x)=\frac{1}{2}" .

"\left | f_{n}(x)- f(x) \right |=\left |\frac{n}{x+2n}-\frac{1}{2} \right | =\left | \frac{2n-x-2n}{x+2n} \right | =\frac{ x}{x+2n}" .

"\frac{x}{x+2n}\leq\frac{k}{2n}" for all "x\in[0,k]" because "x+2n\geq2n" . Therefore, "0\leq\left | f_{n}(x)- f(x) \right | \leq\frac{k}{2n}" and

"0\leq\underset{x\in[0,k] }{sup}\left | f_{n}(x)- f(x) \right | \leq\frac{k}{2n}."

"\lim_{n\rightarrow\infty}\frac{k}{2n }=0" . Hence

"\lim_{ n\rightarrow\infty}\underset{x\in[0,k] }{sup}\left | f_{n}(x)- f(x) \right | =0"

Thus, by the definition, the sequence converges uniformly on "[0,k]."